OpenStudy (anonymous):

what is 12.5 divided by o.25 show ur work

OpenStudy (anonymous):

\[ \frac{12.5}{0.25} = \frac{12.5}{\frac{1}{4}} = 4 *12.5 = 50\]:-)

OpenStudy (kymber):